Question 4 [16 Marks]


Research has shown that about 2% of all products produced by BVH Company are defective. A quality


inspector has just sampled 300 products from this company. Using the Poisson distribution, what is the


probability that:


(a) exactly 12 of them are defective? [4]


(b) at least 8 of them are defective?

1
Expert's answer
2021-11-04T08:43:27-0400

We need to establish the relationship between binomial distribution and Poisson distribution because finding probabilities for number of defectives requires the use of binomial distribution.

Here,

"n=300"

"p=2\%=2/100=0.02"

Relationship between these two distributions is that parameter "\lambda" for a Poisson distribution is the mean of Binomial distribution. This relationship can be written as,

"\lambda=np" where "np" id the mean of a Binomial random varible.

Express by "X" the number of defectives produced by BVH company then, "X\sim Poisson(\lambda)" and its probability mass function is given as,

"p(X=x)=e^{(-\lambda)}(\lambda)^x/{x}!"

The value of "\lambda" is given as,

"\lambda=n*p=0.02*300=6"


a.

Probability of exactly 12 defectives is written as,

"p(X=12)=e^{-6}(6)^{12}\/12!"

By running this command in "R",

"dpois(12,6)= 0.01126448"

Therefore, the probability that there are exactly 12 defectives is 0.01126448


b.

The probability that there are at least 8 defectives is given as,

"p(X\geqslant8)=1-p(X\lt8)=1-\displaystyle\sum^7_{x=0}(e^{-6}6^x)/x!"

Again, we use run the following commands in "R"

"p(X\geqslant8)=1-ppois(7,6)=1-0.7439798= 0.2560202"

Therefore, the probability that there are at least 8 defectives is 0.2560202